OpenStudy (tanya123):

3x2 + 4x2 = 35 as u answer pls explain the method...

OpenStudy (anonymous):

\[ 3x^2 + 4x^2 = 35\] ADD COMMON TERMS \[7x^2 = 35 \] SUBTRACT 35 FROM BOTH SIDES \[7x^2 - 35 = 0\] FACTOR OUR A 7 \[7(x^2 - 5) = 0\] ARE YOU UNDERSTANDING ME SO FAR?

OpenStudy (tanya123):

can u pls tell me where did the 7 come from (ADD COMMON TERMS 7x^2 = 35 )

OpenStudy (anonymous):

okaii do you know had to add? since the 3^2 AND 4x^2 are common terms just add \[3x^2 + 4x^2 = 7x^2\]

OpenStudy (anonymous):

*how

OpenStudy (anonymous):

@tanya123 what are you supposed to do? -FACTOR? -find the solutions?

OpenStudy (tanya123):

okay now i understand where u got the 7 from ...... but what is this sign "^" mean (its not in my question)

OpenStudy (anonymous):

The ^ means "to the"' is the same as writing exponents or those little 2's in the right hand corner.

OpenStudy (tanya123):

ohh.... thank you soo much for helping !!!!!!!